Berlington Inc. expects to earn $18 million per year in perpetuity if it does not undertake any new projects. The firm has an opportunity to invest $10 million today and $10 million in one year in a new project. The new investment will generate annual earnings of $7.5 million in perpetuity, beginning two years from today. The firm has 10 million shares of common stock outstanding, and the required rate of return on the stock is 12 percent.

a) What is the price of a share of stock if the firm does not undertake the new investment?

b) What is the value of the investment?

c) What is the per share stock price if the firm undertakes the investment?
Answer #1

1.
=Earnings/Number of shares*1/required return
=18/10*1/12%=15

2.
=-Present value of outflows+Present value of inflows
=-10-10/1.12+7.5/1.12^2*1/(1-1/1.12)=36.875 million

3.
=Price without investment+NPV of investment/number of shares
=18/10*1/12%+(-10-10/1.12+7.5/1.12^2*1/(1-1/1.12))/10=18.6875